### v1

# INSIGHT: Universal Neural Simulator for Analog Circuits Harnessing Autoregressive Transformers

Abstract.
Analog front-end design heavily relies on specialized human expertise and costly trial-and-error simulations, which motivated many prior works on analog design automation. However, efficient and effective exploration of the vast and complex design space remains constrained by the time-consuming nature of CPU-based SPICE simulations, making effective design automation a challenging endeavor. In this paper, we introduce INSIGHT, a GPU-powered, technology-independent, effective universal neural simulator in the analog front-end design automation loop. INSIGHT accurately predicts the performance metrics of analog circuits across various technology nodes, significantly reducing inference time. Notably, its autoregressive capabilities enable INSIGHT to accurately predict simulation-costly critical transient specifications leveraging less expensive performance metric information. The low cost and high fidelity feature make INSIGHT a good substitute for standard simulators in analog front-end optimization frameworks. INSIGHT is compatible with any optimization framework, facilitating enhanced design space exploration for sample efficiency through sophisticated offline learning and adaptation techniques. Our experiments demonstrate that INSIGHT-M, a model-based batch reinforcement learning framework that leverages INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits. To the best of our knowledge, this marks the first use of autoregressive transformers in analog front-end design.
1. Introduction
Analog circuits vary significantly across and within types, each characterized by distinct performance metrics, making analog front-end design challenging. The heavy reliance on human designer expertise and numerous rounds of costly trial-and-error simulations make the process labor-intensive. The consequent prolonged design cycles and increased time to market have motivated extensive prior research towards its automation.
Various knowledge-based and optimization-based approaches have been proposed to automate analog front-end design, addressing circuit topology selection and device sizing. Knowledge-based methods, such as  1, 2, 3, 4, 5, are time-consuming and lack scalability. Optimization-based approaches, on the other hand, can be split into two categories: equation-based and simulation-based. Equation-based methods like  6, 7, 8, 9, 10, 11, 12 are highly inaccurate on modern technologies and lack scalability, while the simulation-based ones rely on expensive simulations. Evolutionary algorithms like Particle swarm optimization (PSO)  13, advanced differential evolution  14, and genetic algorithms  15, 16 require extensive SPICE simulations (sample inefficient). Bayesian optimization  17, 18 suffers from scalability issues arising from computational demands associated with model updates. Deep Reinforcement Learning (DRL) methods like  19, 20, 21, 22, 23, 24 demonstrate robustness but involve numerous simulator interactions. Modern algorithms  25, 26, 27 use black-box optimization and Model-Based Reinforcement Learning (RL) for device sizing. While they demonstrate improved sample efficiency, they still require many costly simulations. Given the vast and complex design space and the intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance. As a result, data-driven approaches have emerged as a promising solution, representing a recent trend in the field 28, 29.
Initially developed for language tasks 30, 31, transformer architectures 32 have been successfully extended to a diverse array of fields that involve recognizing complex patterns in the data, including vision 33, graph theory 34, and reinforcement learning 35, 36. This expansion was driven by the insight that many problems can be modeled as autoregressive sequence generation tasks. Autoregressive approaches significantly enhance the accuracy of output predictions, ensuring coherent, contextually aware outputs, compared to non-autoregressive approaches, with a minor trade-off in inference speed  37. Transformers have outperformed traditional models like convolutional 38, recurrent 39, and graph neural networks 40, demonstrating superior capabilities in handling complex data. Moreover, their ability to develop foundation models that learn generalized features from data, which can then be fine-tuned for specific tasks with exceptional performance, exemplified by GPT-3 30, has set new trends in AI. However, despite their ability to handle varying input and output sequence lengths, their application for building performance prediction models for analog circuits with diverse circuit parameters and specifications lengths remains unexplored.
In this paper, we introduce INSIGHT, a novel GPU-powered technology-independent universal neural circuit simulator that accurately predicts performance on previously unseen parameters across various circuits and different technologies. Leveraging transformers’ inherent strengths, such as high adaptability and computational efficiency, INSIGHT consistently demonstrates high R2 scores and minimal test losses. Its autoregressive capabilities facilitate accurate prediction of simulation-costly critical transient specifications leveraging performance data from less expensive simulations. These features position INSIGHT as a cost-effective, high-fidelity alternative to standard circuit simulators in analog front-end optimization frameworks. INSIGHT is highly scalable and can seamlessly integrate into existing optimization workflows, facilitating computationally economical and constraint-free design space exploration. This integration significantly enhances traditional EDA optimization schemes, which often face bottlenecks due to computational constraints arising from the compute-intensive nature of simulations required at each design cycle iteration. Additionally, transformers’ robust transfer learning capabilities facilitate INSIGHT’s smooth adaptation to evolving technology landscapes while minimizing data requirements. We integrate INSIGHT within a Model-based Batch RL framework (INSIGHT-M) for circuit sizing and observe very high sample efficiencies for different circuits, underscoring the transformative capabilities of transformers in Analog Front-end Design Automation. To our knowledge, this marks the first application of transformers in Analog Front-end Design Automation.
We have validated our concept on different circuits covering different circuit types including operational amplifiers, trans-impedance amplifiers, comparators, and level shifters across different technology nodes.
The key contributions of this work include:
  • We introduce a novel approach to performance prediction for analog circuits by conceptualizing it as a sequence modeling problem. Inspired by recent advancements in natural language processing, we adapt transformer architectures, which are pivotal to modern language models, to decode intricate behaviors in circuits, thereby enhancing predictive accuracy and efficiency.
  • We propose INSIGHT, an innovative GPU-accelerated technology-independent universal neural circuit simulator. INSIGHT is designed to predict performance across diverse circuits and parameters, offering a versatile tool that significantly improves prediction speed and accuracy.
  • We have rigorously validated our approach across six different circuits across multiple technology nodes. INSIGHT consistently achieves high R2 scores and minimal loss scores across circuits and technology nodes. Its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information, thus demonstrating its capability as a high-fidelity substitute for standard simulators. By embedding INSIGHT within a Model-based Batch RL framework, INSIGHT-M, for circuit sizing optimization, we achieved exceptional sample efficiencies, with at least 50×50$\times$50 × improvement across circuits, thus underscoring INSIGHT’s potential to revolutionize diverse optimization frameworks.
  • To the best of our knowledge, this is the first application of transformer models in analog front-end design automation, paving the way for leveraging recent advancements in language models to enhance circuit design and optimization.
The rest of the paper is organized as follows. Section 2 describes the background and motivation; Section 3 explains the detailed implementation; Section 4 demonstrates the results; Section 5 concludes the paper.
2. Preliminaries
In this section, we delve into the fundamentals of analog circuit physics (Section  2.1) and explore transformer architectures (Section  2.2), discussing our rationale for choosing them. We also highlight the definition of a key design quality metric in Section  2.3, called Figure of Merit (FOM), which will guide our sizing flow.
2.1. Analog Circuit Physics
In analog circuits, performance metrics and parameters are highly interdependent. By leveraging these interdependencies, we can build effective prediction models. For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, including unity-gain bandwidth (UGBW) and phase margin (PM). Similarly, combining all these metrics alongside parameter information provides crucial insights into the circuit’s potential transient behavior. Thus, integrating partial performance information alongside parameter information can significantly enhance prediction accuracies for the other performance specifications, including critical transient specifications such as slew rate and settling time, which are typically costly and slow to simulate. This insight motivates the adoption of autoregressive models for precise performance prediction. These models leverage partial performance information along with parameter information to provide a richer context for predicting the remaining metrics, aiming to build a high-fidelity substitute for standard simulators. Furthermore, given the transferable nature of circuit physics across technology nodes, understanding performance interdependencies within one node provides valuable insights applicable to other nodes. Consequently, a model exhibiting high predictive accuracy for circuit performance on one technology node will likely achieve similar success on subsequent nodes and be easily transferable across advanced technology nodes. These insights, combined with the adaptability and transferability of transformer architectures, motivate the use of autoregressive transformers to build a universal neural simulator that would work efficiently across technologies.
2.2. Transformer Architectures
Transformers  32 constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond. These models, foundational for various state-of-the-art language models, stand out due to their self-attention mechanism, effectively capturing dependencies between tokens in a sequence, regardless of distance. The self-attention mechanism evaluates dependencies within a sequence by computing a weighted sum of value vectors. These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors. Additionally, unlike traditional RNNs 41, transformers process data in parallel, significantly improving efficiency and scalability.
While the original Transformer architecture 32 consists of stacked layers of encoders and decoders, decoder-only models like GPT 42 have recently demonstrated notable success in sequence data generation. Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks. Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence. This architectural choice is optimized for autoregressive tasks, where the model iteratively predicts the subsequent token by leveraging all previously generated tokens along with the input sequence as context.
From a probabilistic perspective, in the context of autoregressive generation, transformers model the joint probability of a sequence of tokens ($x_{1}$,$x_{2}$,…,$x_{N}$) by decomposing it into a product of conditional probabilities:
(1) $P(x_{1},x_{2},\dots,x_{N})=\prod_{i=1}^{N}P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$
Each term $P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$, which quantifies the likelihood of observing the token $x_{i}$, given the sequence of all preceding tokens ($x_{1},x_{2},\dots,x_{i-1}$), is estimated using the self-attention mechanism. The mechanism’s ability to capture long-range dependencies and contextual relationships within data, along with its adaptability to variable sequence lengths, makes Transformers ideal for handling complex sequences with widely varying input sizes. Drawing inspiration from this probabilistic formulation, we aim to enhance analog circuit performance prediction by strategically integrating partial performance data with parameter information, providing richer context as detailed in the following section.
2.3. Figure of Merit
Analog circuit sizing involves determining the optimal device sizes to meet specified performance criteria. Formally, given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  25, 43, 26, to evaluate quality of a design. This is of the following form:
(2) $\text{FoM}(\mathbf{x})=w_{0}\cdot f_{0}(\mathbf{x})+\sum_{i=1}^{m}\min\left(1,\max(0,w_{i}\cdot f_{i}(\mathbf{x}))\right)$
where $w_{i}$ is the weighting factor, a $\max(\cdot)$ function is used for equating designs after constraints are met, and $\min(\cdot)$ is used to prevent a single constraint violation from dominating the FoM value.
3. Algorithms
In this section, we introduce our proposed frameworks. We begin by detailing the construction of our proposed universal neural simulator for analog circuits, INSIGHT. Figure 1 illustrates how INSIGHT conceptualizes the prediction of performance metrics as an autoregressive sequence generation task, with Section 3.1 providing the probabilistic rationale for this approach. Section 3.2 then offers a detailed discussion of our analog circuit sizing optimization algorithm leveraging INSIGHT. As illustrated in Figure 2, the analog circuit sizing optimization algorithm for a given topology consists of two primary phases. The initial phase, discussed in Section 3.2.1, consists of a one-time effort for simulation-based performance dataset generation and training of INSIGHT. The second phase, discussed in Section 3.3, involves the integration of the one-time trained INSIGHT into a Model-Based Batch RL framework, termed INSIGHT-M, to perform automated analog sizing optimization for previously unseen user-input target specifications.
3.1. INSIGHT
Formally, for a given analog circuit topology, given a sequence of design parameters ($x_{1}$ to $x_{N}$), such as width-to-length ratios (W/L), our goal is to predict the corresponding performance specifications ($y_{1}$ to $y_{M}$) accurately. This predictive task is mathematically framed as maximizing the conditional probability $P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})$, which can be expressed as follows:
(3) $\begin{split}P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})=\\ \prod_{i=1}^{M}P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})\end{split}$
This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
(4) $\text{NLL}=-\sum_{i=1}^{M}\log P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})$
This formulation mirrors the contextual approach used in autoregressive sequence generation tasks, such as those involving decoder-based transformers designed for enhanced sequence analysis. In analog circuits, where performance metrics are inherently interdependent, this formulation attempts to leverage these intricate relationships to predict the performance metrics for previously unseen input parameters accurately. Figure 1 presents a conceptual diagram illustrating how INSIGHT implements autoregressive modeling to predict performance metrics. The self-attention mechanism within the Transformer model enables it to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$. This capability is crucial for uncovering insights into the physics of the circuit, as it allows the model to learn and utilize the dependencies of one performance metric on another along with parameter information. Table 1 presents INSIGHT’s architecture details.
Figure 1. Autoregressive circuit performance prediction in INSIGHT
3.2. Analog Circuit Sizing Optimization
Figure 2. Automatic Analog Sizing Framework
Figure 2 showcases our sizing optimization algorithm. It comprises of two main sections: INSIGHT Pre-Training Phase (Section  3.2.1) and INSIGHT-M (Section  3.3). The core of RL in this paper is PPO 44. We basically use the hyper-parameters therein. Note that other RL algorithms can be employed with INSIGHT.
Table 1. INSIGHT Architecture
INSIGHT Parameter Value
  Network Architecture Decoder-Based Transformer
  Dim. of a model 76
  # of Heads 4
  # of Decoder Layers 3
  Activation GeLU
Table 2. Neural Simulator Architecture in CRONuS
FC-Ensemble Parameter Value
 Network Architecture Fully-Connected
# of Hidden Units 200-200-200-200-200
# of Ensemble Networks 7
Activation ReLU
3.2.1   INSIGHT Pre-Training
As depicted in the left part of Figure 2, the pre-training phase involves a one-time effort to generate a simulation-based performance dataset, followed by a one-time training of our proposed neural simulator, INSIGHT. The data collection involves gathering a random dataset of Gaussian-sampled design parameters and their corresponding performance metrics derived from SPICE simulations. To address computational demands, the initial dataset size is defined as a hyperparameter, allowing for adjustments tailored to specific use cases.
This flexibility helps balance the resource usage during the initial data collection with the subsequent sample efficiency following a user query. This process is followed by the one-time pre-training of our proposed neural simulator, INSIGHT. By employing transformer technology, INSIGHT effectively utilizes transfer learning, thereby reducing dataset requirements across different technologies.
3.3. INSIGHT-M
The pre-trained INSIGHT is subsequently integrated into a Model-Based Batch RL framework for circuit sizing, referred to as INSIGHT-M. As shown in the right part of Figure 2, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance target specifications, including previously unseen ones. In our experiments, we utilized a Proximal Policy Optimization (PPO)-based RL agent 44, a widely used approach in Reinforcement Learning. PPO has been extensively utilized in analog design optimization, as in  45, 24. PPO operates by executing multiple trajectory rollouts in parallel, followed by their subsequent aggregation to update the policy under certain constraints. Although PPO’s use of multiple trajectories under safe constraints enhances the robustness and safety of the learning process, it tends to compromise sample efficiency. The INSIGHT-M framework addresses this drawback by enabling the RL agents to be pre-trained within the close-to-real synthetic neural simulator environment, INSIGHT, until they achieve satisfactory performance. Once trained, these agents are deployed in the real-world simulator environment using a batch RL strategy. This process involves periodic performance evaluations in the real environment. If the agents perform adequately, their deployment persists. If not, new real-world data is acquired to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated. This iterative refinement process optimizes the balance between safety, robustness, and efficiency.
Central to the efficiency of INSIGHT-M is the training of a robust RL agent within a synthetic environment that closely mimics real-world conditions. The high fidelity of INSIGHT ensures that the agent’s learning and subsequent behaviors are closely aligned with real-world design tasks, thereby minimizing trial and error and improving sample efficiency. Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, enhances the agent’s adaptability to real-world environments. Furthermore, the framework’s capability for iterative refinement effectively balances exploration and exploitation, achieving high sample efficiency.
4. Experiments
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e) (f)
Figure 3. Benchmark Circuits: (a) Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, (b) Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, (c) Two-Stage Transimpedance Amplifier, (d) Three-Stage Transimpedance Amplifier, (e) Comparator, (f) Level Shifter.
Table 3. Performance Comparison Across Different Circuits: INSIGHT vs Neural Simulator in CRONuS
Circuits (45nm) R2 Scores
Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¿0.99 0.65 ¿0.99 0.62
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¿0.99 0.73 ¿0.99 0.77
Two-Stage Transimpedance Amplifier ¿0.99 0.81 ¿0.99 0.78
Three-Stage Transimpedance Amplifier ¿0.99 0.91 ¿0.99 0.88
Comparator 0.70 0.60 0.97 0.61
Level Shifter ¿0.99 ¡0.00 0.91 0.07
Circuits (45nm) MSE
Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¡0.001 0.01 ¡0.001 0.01
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¡0.001 0.05 ¡0.001 0.04
Two-Stage Transimpedance Amplifier ¡0.001 0.05 ¡0.001 0.05
Three-Stage Transimpedance Amplifier ¡0.001 0.01 ¡0.001 0.01
Comparator 0.010 0.01 ¡0.001 0.01
Level Shifter ¡0.001 0.03 ¡0.001 0.05
Table 4. INSIGHT Performance Across Different Technologies
Technology 45nm 130nm 180nm
Circuit Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Two-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Three-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Table 5. Performance Comparison of INSIGHT-M (with PPO) vs AutoCkt for 45nm case
Algorithm INSIGHT-M (with PPO) AutoCkt
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1,500 steps 14,100 steps
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1,500 steps 9,300 steps
Two-Stage Transimpedance Amplifier 300 steps 11,100 steps
Three-Stage Transimpedance Amplifier 600 steps 8,700 steps
Comparator 1,200 steps 8,400 steps
4.1. Benchmark Tasks and Algorithms
In this paper, we utilized a GPU machine with x8 NVIDIA P100 GPUs under Intel(R) Xeon(R) CPU E5-2698 V4 2.220GHz for training and inference. We used the following benchmark circuits to evaluate INSIGHT and to validate our approach: Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, Two-Stage Transimpedance Amplifier, Three-Stage Transimpedance Amplifier, Comparator, and Level Shifter. Figure 3 illustrates the detailed topology of the benchmark circuits. Table 6 presents the performance specifications for each circuit we focussed on for this paper. We have validated our concept using BSIM 45nm, Skywater 130nm, BSIM 180nm, and FreePDK 45nm technology nodes.
4.2. INSIGHT Performance
4.2.1   Prediction Accuracy
Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  27, which serves as the baseline for our comparison with INSIGHT. Table 3 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on the BSIM 45nm technology node. As demonstrated in the table, our analysis for the comparator and level shifter cases involved a combination of DC and simulation-costly transient specifications. Specifically for the level shifter, leveraging its autoregressive capabilities, INSIGHT enhances the predictability of simulation-intensive transient specifications such as Avg. Delay and Delay Balance by efficiently utilizing economical metric information like DC Power and Ratio. The table further details INSIGHT’s performance across various training and testing data sizes, indicating robust performance for dataset sizes close to 1500-1800. Given that simulation data collection is a one-time effort, this investment can prove cost-effective if a topology is to be optimized multiple times. In our optimization framework, we define the initial dataset size as a hyperparameter to efficiently balance computational resource utilization with subsequent sample efficiency, allowing for adjustments tailored to specific use cases. Table 4 validates INSIGHT’s effectiveness across 45nm, 130nm, and 180nm technology nodes, reinforcing its universal applicability. Our experiments showed that simple transfer learning or knowledge-based domain adaptation techniques can reduce initial dataset requirements.
4.2.2   Inference Speed
We evaluated INSIGHT’s performance prediction inference time across our benchmark circuits. When simultaneously processing a batch of 1,000 combinations of input parameters, the average inference time across circuits was $115.21\pm 12.01\mu s$. This result was derived from 100 trials per circuit on our setup.
4.3. INSIGHT-M Performance
We evaluate the performance of INSIGHT-M against AutoCkt 24, an established online learning-based sizing algorithm based on Proximal Policy Optimization (PPO), using default hyperparameter settings. Table 5 highlights the performance comparison. Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance. For these results, the INSIGHT model used within INSIGHT-M was pre-trained on a minimal initial dataset of 300 samples, showcasing the efficacy of our neural simulator INSIGHT in optimization frameworks.
Table 6. Performance Specifications for each circuit
Circuit Performance Specifications Width (Capacitor) Range ($\mu$m ($pF$))
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair $I_{Q}$ (mA) DC Gain (dB) UGBW (MHz) Phase Margin (degrees) [1.0e-1, 3.0e1]
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair [1.0e-1, 3.0e1]
Two-Stage Transimpedance Amplifier [1.0e-1, 1.0e2]
Three-Stage Transimpedance Amplifier [1.0e-2, 3.0e1]
Comparator DC Power (W) Avg. Delay (sec) [1.0e-1, 3.0e1]
Level Shifter DC Power (W) Ratio Avg. Delay (sec) Delay Balance (sec) [4.6e-1, 1.0e2]
5. Conclusion
This paper proposes INSIGHT, a novel, GPU-powered, technology-independent, effective universal neural simulator in the analog front-end automation loop. Leveraging self-attention to understand underlying inherent parameter-performance interdependencies, INSIGHT accurately predicts performance across analog circuits and their corresponding metrics over various technology nodes, achieving inference times within microseconds. Notably, its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information. Due to its low cost and high fidelity, INSIGHT serves as an effective substitute for standard simulators in analog front-end optimization frameworks, facilitating effective and efficient design space exploration and improved sample efficiency. Our Model-Based Batch RL framework, INSIGHT-M, which utilizes INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits. We rigorously validated our approach on diverse circuit types across technology nodes. To our knowledge, this marks the first application of autoregressive transformers in analog front-end design automation. We plan to extend this technology to address more complex analog design challenges, including mismatch analysis and design migration. We aim to enhance our adaptation techniques to mitigate synthetic to real-world adaptation errors, thereby strengthening the robustness and reliability of our models under dynamic conditions. This innovative application of autoregressive transformers opens new avenues for leveraging advanced NLP techniques, including grammatical models and language-based inversion techniques in circuit modeling and optimization, inspiring further innovation in the field.